# Unexpected but informative: What fixation-related potentials tell us about the processing of ambiguous program code

### v1

@@ Abstract
As software pervades more and more areas of our professional and personal lives, there is an ever-increasing need to maintain software, and for programmers to be able to efficiently write and understand program code. In the first study of its kind, we analyze fixation-related potentials (FRPs) to explore the online processing of program code patterns that are ambiguous to programmers, but not the computer (so-called atoms of confusion), and their underlying neurocognitive mechanisms in an ecologically valid setting. Relative to unambiguous counterparts in program code, atoms of confusion elicit a late frontal positivity with a duration of about 400 to 700 ms after first looking at the atom of confusion. As the frontal positivity shows high resemblance with an event-related potential (ERP) component found during natural language processing that is elicited by unexpected but plausible words in sentence context, we take these data to suggest that the brain engages similar neurocognitive mechanisms in response to unexpected and informative inputs in program code and in natural language. In both domains, these inputs lead to an update of a comprehender’s situation model that is essential for information extraction from a quickly unfolding input.
@@
program comprehension, natural language comprehension, atoms of confusion, EEG, fixation-related potentials, event-related potentials, late frontal positivity1 Introduction
@@
The main goal of the present study is to explore the processing of ambiguous and therefore potentially confusing elements in program code and their underlying neurocognitive mechanisms in an ecologically valid setting. Recently, various studies identified a set of small, recurring patterns [7] prevalent in program code [8, 9, 10], the so-called atoms of confusion. An atom of confusion is a syntactically correct code pattern that is unambiguous for the computer to execute, but, due to its unintuitive structure, is ambiguous to programmers, who encounter difficulties in understanding the code and might even misinterpret it [7, 2].
lstlisting ⬇ 1int V1 = 2; 2int R = 3 + V1; 3V1++; Listing 2: Equivalent unambiguous code without atom of confusion (the increment is extracted as a separate statement)
In Figure 1, we contrast an example of an ambiguous code snippet that contains an atom of confusion with an unambiguous code snippet that contains functionally-equivalent code without the atom of confusion. The atom of confusion is located in Line 2 of the ambiguous snippet displayed in Listing 1, which combines an addition (+) and a post-increment operator (V1++, equivalent to V1 = V1 + 1) within one statement. The difficulty that programmers may encounter is that it might be unclear whether the post-increment operator is executed before assigning the value of the sum to variable R or after, leading either to the incorrect result of R = 6 or the correct result of R = 5. In the unambiguous counterpart shown in Listing 2, the increment is extracted to a separate statement in Line 3. This eliminates any ambiguity regarding the order of execution, and the output of the variable R can be unambiguously calculated as 5. It is important to note that the ambiguous and unambiguous snippet conditions are both syntactically correct and semantically equivalent, but they differ in their chance that a human programmer may misunderstand their semantics [7] due to the ambiguity arising from the atom of confusion. Several behavioral studies across different programming languages have shown that atoms of confusion require more time to process [3, 7], are more effortful to comprehend [15, 13], and lead to more errors [7, 16, 1].
@@
Despite the insights on atoms of confusion obtained by behavioral experiments, the underlying mechanisms of the phenomenon are unclear. One missing element is our limited understanding of which neurocognitive mechanisms exactly are engaged in response to the ambiguous program code that is causing the difficulty. Identifying these mechanisms would allow avoiding misconceptions in program code and eliminating problematic programming patterns, with implications for designing programming languages, assistants, and curricula.
@@
The premise of our work is that our understanding of how humans process natural language could be the key to interpreting the processing of atoms of confusion and discovering the underlying neurocognitive mechanisms. Indeed, parallels between programming languages and natural languages have been observed in several studies, hinting at an overlap in activated brain areas between natural language (e.g., Brodmann Areas 21, 44, and 47) and program comprehension [17, 18, 19, 20, 21]. In general, considering analogies with what is already known about human language processing and the processing of program code might prove to be a very fruitful avenue.
@@
Or should we rather think of atoms of confusion in terms of (2a) syntactic reanalysis similar to what is typically observed in garden path sentences, such as The horse raced past the barn fell, which, like an atom of confusion, in fact only has a single correct interpretation, which is however difficult to work out for a human, as there might be uncertainty about the function of the verb raced. It can either be analyzed as a main verb of the sentence or alternatively as a past participle of a reduced relative clause The horse (that was) raced past the barn fell (which is similar to V1++ in our code example, where there is uncertainty about the program execution order). In this case, we should expect to see a late posterior positivity (P600) [28, 29]. Besides garden path sentences that involve syntactic reanalysis, the P600 is generally observed in sentences with syntactic errors, such as The child throw the toy, but also in sentences with syntactic ambiguities, as in The woman persuaded to answer the door relative to the syntactically unambiguous sentence The woman struggled to prepare the meal [29, 30]. In these cases, the P600 is typically interpreted as a correlate of a syntactic revision (e.g., passivizing the verb persuaded and attaching it to a reduced relative clause), but it can also be evoked by semantic violations (see Paczynski and Kuperberg [31] for an example).
@@
Currently, it is unknown whether the ambiguity induced by atoms of confusion as perceived by programmers results (1) from semantic integration of unexpected and challenging content (in which case we would expect to see an N400), (2) from syntactic ambiguity (see below), or (3) from something else entirely (producing an unexpected ERP signal). For the syntactic ambiguity, the ambiguity may arise from (2a) syntactic reanalysis (in which case we would expect to see a late posterior positivity) or from (2b) integration of unexpected but plausible content into the situation model of the preceding context (in which case we would expect to see a late frontal positivity). The premise of our work is that building on natural-language studies would help us to understand the source and processing of ambiguity, illuminating the intricacies of program comprehension. This, in turn, would be invaluable to better support programmers by dedicated programming languages and curricula. For example, if atoms of confusion evoke a perceived syntactic ambiguity, this could be avoided by better design choices of programming language designers; if they rather cause problems with semantic integration, this can be captured in coding guidelines and dedicated training procedures in programming curricula. If we observe an entirely unexpected ERP signal, this could indicate that the ambiguity arising from the presence of an atom of confusion has no equivalent in natural-language comprehension, or that there is no unified cause for atoms of confusion. Thus, independent of the concrete results, it is highly relevant for the field of program comprehension to understand the electrophysiological signals and hence the underlying mechanisms of processing atoms of confusion. Looking at a broader picture, if we observe more and more analogies between processing program code and natural language text, this suggests that our premise of a methodological transfer between the two disciplines is worthwhile.
@@
Since it is unclear which ERP component relates to processing ambiguous code, we employ a data-driven approach to detect differences between ambiguous and corresponding unambiguous code without targeting a hypothesized component, as research on program comprehension and its underlying neurocognitive mechanisms is still in its infancy. Contrary to research on natural language comprehension, which looks back on several decades of behavioral, neurophysiological, and brain imaging studies that have unraveled a variety of processing aspects and representational structures [35], studying program comprehension faces unique challenges, such as non-linear reading flow (e.g., program statements are processed not in lexical order but along control and data flow of possible program executions) [36, 37, 38, 39]. So far, only one study on program comprehension investigated ERP responses to syntactic and semantic errors in program code: Using a token-wise presentation of code, Kuo and Prat [40] identified P600 and N400 components for syntactically and semantically invalid program code, respectively. However, this setup of a token-wise presentation, while usually applied in reading studies with natural language materials [41], is too artificial and highly problematic for the analysis of program code: First, it enforces an unnatural linear reading flow. Second, with token-wise presentation mode, each token is presented with a uniform duration, which allows the programmer only a cursory glance at the code but does not account for calculations that can strongly vary in complexity between tokens within a code snippet. Both aspects strongly compromise the construct and ecological validity of neurocognitive studies on program comprehension.
@@
To uncover which neurocognitive mechanisms exactly are engaged in response to processing ambiguous program code, we conducted an FRP experiment involving participants. We carefully adjusted the existing stimuli from previous work on atoms of confusion in code written in the Java programming language [1] to optimize them for an FRP setup. In particular, each snippet was presented in a condition with and without an atom of confusion (across participants). Each pair of the snippet condition had a comparable number and length of code lines. Furthermore, to select suitable fixations for FRP analysis, we ensured that each snippet contained a single, clearly defined area of interest that contains either the atom of confusion (ambiguous snippets) or an equivalent without atom (unambiguous snippets) (see Section 4, for details). The experiment consisted of three blocks, as illustrated in Figure 2, in which we presented snippets (half ambiguous and half unambiguous). We asked the participants to comprehend the snippets as fast and accurately as possible in order to be able to report the output. During the experiment, we collected comprehension time, answer correctness, and subjective difficulty ratings to verify that our ambiguous snippets were generally more difficult to process on the behavioral level than unambiguous snippets. We collected EEG and eye-tracking data to calculate FRPs by mapping fixations to the areas of interest within the code snippets. We took special care to identify and correct eye movements and muscular artifacts by means of an independent component analysis (see Section 4, for further details). We synchronized the fixations with the EEG data stream to calculate FRPs. On the obtained FRPs, we tested for significant amplitude differences between ambiguous and unambiguous code using pairwise cluster permutation tests.
@@ 2 Results
We report on the first-ever FRP analysis of program comprehension by measuring and pinpointing the processing of ambiguous elements and its neural underpinning during program comprehension in an ecologically valid setting. Our results show that ambiguous code elicits a larger frontal positivity between and after first fixating the relevant area of the atom of confusion compared to fixating unambiguous code (without atom). We relate the identified component to the late frontal positivity occurring in natural language processing for unexpected but plausible input which elicits an update of the situation model.
@@ 2.1 Behavioral results
In line with prior work, our behavioral results based on general linear mixed-effects regression (GLMER) models show that the ambiguous snippets took longer to comprehend, had fewer correct answers, and were rated less often as easy to understand, as displayed in the right part of Table 1. The full analysis results including all fixed and random effects are provided in the supplementary information, Appendix C.
@@ 2.2 Fixation-related potentials analysis
In a first step, we performed an ERP analysis of the EEG time-locked to the onset of the snippets with the goal of testing whether, without the eye-tracking information, ERPs differ between ambiguous and unambiguous code. Consistent with our assumption that ERPs cannot reveal meaningful differences when presenting a code snippet as a whole, we found no significant ERP differences between the two conditions (cf. supplementary information, Appendix B).
@@
Our primary goal was to determine the neurocognitive processes triggered by the presence of atoms of confusion in program code and to relate those processes to ERP components typically identified during natural language comprehension. In fact, there is a large delay and large variance in the point in time when participants first fixate the critical region after onset of the snippet: the first fixation on the critical region happens, on average, after stimulus onset in the unambiguous condition, and in the ambiguous condition. The stimulus-onset aligned ERPs hence cannot capture this first point in time at which information from the critical region is presumably processed. Fixation data show that on the first pass, participants fixated the critical region for in the unambiguous condition and in the ambiguous condition. Therefore, we compared FRPs time-locked to the first fixation in the area of the atom of confusion and its counterpart without atom, spanning an epoch from preceding the start of the fixation until thereafter. For some trials, this epoch could not be extracted due to either bad quality in the fixation data, no fixation in the area of interest, artifacts in the EEG signal, or a too short time period until the end of the snippet viewing phase. We excluded these trials from the subsequent analysis ( of ). Then, we calculated the participants’ FRP averages for both conditions across all trials. The grand average FRP data across all participants for electrodes at frontal, central, and parietal recording sites are presented in Figure 3a.
@@
As illustrated in Figure 3, ambiguous code elicits a sustained frontal positivity relative to unambiguous code. This positivity was most pronounced between and at frontal recording sites (i.e., F3, Fz, F4). At posterior recording sites, the positive amplitude difference between ambiguous and unambiguous code was attenuated in this time interval. Figure 3b shows the difference in amplitude for all electrodes in form of topographical distributions in consecutive time intervals between and . Positive differences between both conditions were also present beyond after fixation onset at some left frontal recording sites (e.g., F7 and F3), though less consistent. We statistically analyzed these effects using a cluster-based paired permutation t-test, which identifies significant clusters of adjacent data points (defined temporally and spatially).
@@
The results from the permutation test with the alternative hypothesis reveal one significant cluster () in the time interval from to (cf. Figure 3c). Although almost all electrodes are considered to be part of the significant cluster (at least, at one time point), the largest temporal extension of significant amplitude differences can be observed at midline frontal and fronto-central recording sites (i.e., Fz, FCz). Figure 3c displays the significant cluster across time and electrodes (surrounding black lines) over the amplitude difference in the grand average. For the alternative hypothesis , we detected no significant clusters.
@@
These results show that the late frontal positivity is significantly greater for ambiguous code than for unambiguous code, indicating that both conditions are processed differently starting at around after onset of the first fixation.

### v2

# Unexpected but informative: What fixation-related potentials tell us about the processing of confusing program code

@@ Abstract
As software pervades more and more areas of our professional and personal lives, there is an ever-increasing need to maintain software, and for programmers to be able to efficiently write and understand program code. In the first study of its kind, we analyze fixation-related potentials (FRPs) to explore the online processing of program code patterns that are confusing to programmers, but not the computer (so-called atoms of confusion), and their underlying neurocognitive mechanisms in an ecologically valid setting. Relative to clean counterparts in program code without an atom of confusion, confusing code elicits a late frontal positivity of about 400 to 700 ms after first looking at the atom of confusion. This frontal positivity resembles an event-related potential (ERP) component found during natural language processing that is elicited by unexpected but plausible words in sentence context. Thus, we suggest that the brain engages similar neurocognitive mechanisms in response to unexpected and informative inputs in program code and in natural language. In both domains, these inputs update a comprehender’s situation model that is essential for information extraction from a quickly unfolding input. Our results have far-reaching implications for programming and pave the way for interdisciplinary collaborations between software engineering and psycholinguistics.
@@
program comprehension, natural language comprehension, atoms of confusion, EEG, fixation-related potentials, late frontal positivity1 Introduction
@@
The main goal of the present study is to explore the processing of confusing elements in program code and their underlying neurocognitive mechanisms in an ecologically valid setting. Recently, various studies identified a set of small, recurring patterns [7] prevalent in program code [8, 9, 10], the so-called atoms of confusion. An atom of confusion is a syntactically correct code pattern that is unambiguous for the computer to execute, but, due to its unintuitive structure, is confusing to programmers, who encounter difficulties in understanding the code and might even misinterpret it [7, 11].
In Figure 1, we contrast an illustrative example of a confusing code snippet (containing an atom of confusion) with a clean code snippet (containing functionally-equivalent code without the atom of confusion). The atom of confusion is located in Lines 7 and 8 of the confusing snippet displayed in Listing 1, in which the level of indentation of the else block subtly influences human understanding. The difficulty that programmers may encounter is that the indentation suggests the else block closes the first if statement, which leads to the incorrect result of
R = 7. In the clean counterpart shown in Listing 2, the indentation aligns with the functionally correct second if statement. This eliminates the risk of confusion regarding the matching if statement, and the output of the variable
R can be straightforwardly determined as 3, as several studies on atoms of confusion confirmed [12, 7, 13]. It is important to note that the confusing and clean snippet conditions are both syntactically correct and semantically equivalent, but they differ in their risk that a human programmer may misunderstand their semantics due to the atom of confusion [7]. While such subtle differences may seem insignificant at first glance, they have caused severe bugs in the past, such as a mismatching indentation in Apple’s iOS operating system rendering its secure network connections vulnerable [14, 15].
In general, several behavioral studies across different programming languages have shown that atoms of confusion require more time to process [12, 7], are more effortful to comprehend [16, 17], and lead to more errors [7, 18, 13].
@@
Despite the insights on atoms of confusion obtained by behavioral experiments, the underlying mechanisms of the phenomenon are unclear. One missing element is our limited understanding of which neurocognitive mechanisms exactly are engaged in response to the confusing program code that is causing the difficulty. Identifying these mechanisms would allow us to more deeply understand the principles behind misconceptions in program code, with far-reaching implications for programming practice. Specifically, this line of research can provide empirical evidence for refining coding conventions, programming language design, and educational practices, ultimately leading to more secure and more reliable software.
@@
The premise of our work is that our understanding of how humans process natural language could be the key to interpreting the processing of atoms of confusion and to discovering the underlying neurocognitive mechanisms. Indeed, parallels between programming languages and natural languages have been observed in several studies, hinting at an overlap in activated brain areas between natural language (e.g., Brodmann Areas 21, 44, and 47) and program comprehension [19, 20, 21, 22, 23]. In general, considering analogies with what is already known about human language processing and the processing of program code might prove to be a very fruitful avenue.
@@
Or should we rather think of atoms of confusion in terms of (2a) syntactic reanalysis similar to what is typically observed in garden path sentences, such as The horse raced past the barn fell, which, like an atom of confusion, in fact only has a single correct interpretation, which is however difficult to work out for a human, as there might be uncertainty about the function of the verb raced. It can either be analyzed as a main verb of the sentence or alternatively as a past participle of a reduced relative clause The horse (that was) raced past the barn fell (which is similar to indentation in our code example, where there is uncertainty about the program execution). In this case, we should expect to see a late posterior positivity (P600) [30, 31]. Besides garden path sentences that involve syntactic reanalysis, the P600 is generally observed in sentences with syntactic errors, such as The child throw the toy, but also in sentences with syntactic ambiguities, as in The woman persuaded to answer the door relative to the syntactically unambiguous sentence The woman struggled to prepare the meal [31, 32]. In these cases, the P600 is typically interpreted as a correlate of a syntactic revision (e.g., passivizing the verb persuaded and attaching it to a reduced relative clause), but it can also be evoked by semantic violations (see Paczynski and Kuperberg [33] for an example).
@@
Currently, it is unknown whether the confusion induced by atoms of confusion as perceived by programmers results (1) from semantic integration of unexpected and challenging content (in which case we would expect to see an N400), (2) from syntactic ambiguity (see below), or (3) from something else entirely (producing an unexpected ERP signal). As for syntactic ambiguity, the confusion may arise from (2a) syntactic reanalysis (in which case we would expect to see a late posterior positivity) or from (2b) integration of unexpected but plausible content into the situation model of the preceding context (in which case we would expect to see a late frontal positivity). The premise of our work is that building on natural-language studies would help us to understand the source and processing of such confusion, illuminating the intricacies of program comprehension. This, in turn, would be invaluable to better support programmers by dedicated programming languages and curricula. For example, if atoms of confusion evoke a perceived syntactic ambiguity, this could be avoided by better design choices of programming language designers (e.g., by requiring explicit brackets around bodies of for, while, and if statements); if they rather cause problems with semantic integration, this can be captured in dedicated training procedures in programming curricula and coding guidelines (e.g., defining clear and consistent rules that instruct programmers on the use of patterns). If we observe an entirely unexpected ERP signal, this could indicate that the confusion arising from the presence of an atom of confusion has no equivalent in natural-language comprehension, or that there is no unified cause for atoms of confusion. Thus, independent of the concrete results, it is highly relevant for the field of program comprehension to understand the electrophysiological signals and hence the underlying mechanisms of processing atoms of confusion. Looking at a broader picture, if we observe more and more analogies between processing program code and natural language text, this suggests that our premise of a methodological transfer between the two disciplines is worthwhile.
@@
Since it is unclear which ERP component relates to processing confusing code, we employ a data-driven approach to detect differences between confusing and corresponding clean code without targeting a hypothesized component, as research on program comprehension and its underlying neurocognitive mechanisms is still in its infancy. Contrary to research on natural language comprehension, which looks back on several decades of behavioral, neurophysiological, and brain imaging studies that have unraveled a variety of processing aspects and representational structures [37], studying program comprehension faces unique challenges, such as non-linear reading flow (e.g., program statements are processed not in lexical order but along control and data flow of possible program executions) [38, 39, 40, 41]. So far, only one study on program comprehension investigated ERP responses to syntactic and semantic errors in program code: Using a token-wise presentation of code, Kuo and Prat [42] identified P600 and N400 components for syntactically and semantically invalid program code, respectively. However, this setup of a token-wise presentation, while usually applied in reading studies with natural language materials [43], is too artificial and highly problematic for the analysis of program code: First, it enforces an unnatural linear reading flow. Second, with token-wise presentation mode, each token is presented with a uniform duration, which allows the programmer only a cursory glance at the code but does not account for calculations that can strongly vary in complexity between tokens within a code snippet. Both aspects strongly compromise the construct and ecological validity of neurocognitive studies on program comprehension.
@@
To uncover which neurocognitive mechanisms exactly are engaged in response to processing confusing program code, we conducted an FRP experiment involving participants. We carefully adjusted the existing stimuli from previous work on atoms of confusion in code written in the Java programming language [13] to optimize them for an FRP setup. In particular, each snippet was presented in a condition with and without an atom of confusion (across participants). Each pair of the snippet condition had a comparable number and length of code lines. Furthermore, to select suitable fixations for FRP analysis, we ensured that each snippet contained a single, clearly defined area of interest that contains either the atom of confusion (confusing snippets) or an equivalent without atom (clean snippets) (see Section 4, for details). The experiment consisted of three blocks, as illustrated in Figure 2, in which we presented snippets (half confusing and half clean). We asked the participants to comprehend the snippets as fast and accurately as possible in order to be able to report the output. During the experiment, we collected comprehension time, answer correctness, and subjective difficulty ratings to verify that our confusing snippets were generally more difficult to process on the behavioral level than clean snippets. We collected EEG and eye-tracking data to calculate FRPs by mapping fixations to the areas of interest within the code snippets. We took special care to identify and correct eye movements and muscular artifacts by means of an independent component analysis (see Section 4, for further details). We synchronized the fixations with the EEG data stream to calculate FRPs. On the obtained FRPs, we tested for significant amplitude differences between confusing and clean code using pairwise cluster permutation tests.
@@ 2 Results
We report on the first-ever FRP analysis of program comprehension by measuring and pinpointing the processing of confusing elements and its neural underpinning during program comprehension in an ecologically valid setting. Our results show that confusing code elicits a larger frontal positivity between and after first fixating the relevant area of the atom of confusion compared to fixating clean code (without an atom). We relate the identified component to the late frontal positivity occurring in natural language processing for unexpected but plausible input which elicits an update of the situation model.
@@ 2.1 Behavioral results
In line with prior work, our behavioral results based on general linear mixed-effects regression (GLMER) models show that the confusing snippets took longer to comprehend, had fewer correct answers, and were rated less often as easy to understand, as displayed in the right part of Table 1. The full analysis results including all fixed and random effects are provided in the supplementary information, Appendix C.
@@ 2.2 Fixation-related potentials analysis
In a first step, we performed an ERP analysis of the EEG time-locked to the onset of the snippets with the goal of testing whether, without the eye-tracking information, ERPs differ between confusing and clean code. Consistent with our assumption that ERPs cannot reveal meaningful differences when presenting a code snippet as a whole, we found no significant ERP differences between the two conditions (cf. supplementary information, Appendix B).
@@
Our primary goal was to determine the neurocognitive processes triggered by the presence of atoms of confusion in program code and to relate those processes to ERP components typically identified during natural language comprehension. In fact, there is a large delay and large variance in the point in time when participants first fixate the critical region after onset of the snippet: the first fixation on the critical region happens, on average, after stimulus onset in the clean condition, and in the confusing condition. The stimulus-onset aligned ERPs hence cannot capture this first point in time at which information from the critical region is presumably processed. Fixation data show that on the first pass, participants fixated the critical region for in the clean condition and in the confusing condition. Therefore, we compared FRPs time-locked to the first fixation in the area of the atom of confusion and its counterpart without atom, spanning an epoch from preceding the start of the fixation until thereafter. For some trials, this epoch could not be extracted due to either bad quality in the fixation data, no fixation in the area of interest, artifacts in the EEG signal, or a too short time period until the end of the snippet viewing phase. We excluded these trials from the subsequent analysis ( of ). Then, we calculated the participants’ FRP averages for both conditions across all trials. The grand average FRP data across all participants for electrodes at frontal, central, and parietal recording sites are presented in Figure 3a.
@@
As illustrated in Figure 3, confusing code elicits a sustained frontal positivity relative to clean code. This positivity was most pronounced between and at frontal recording sites (i.e., F3, Fz, F4). At posterior recording sites, the positive amplitude difference between confusing and clean code was attenuated in this time interval. Figure 3b shows the difference in amplitude for all electrodes in form of topographical distributions in consecutive time intervals between and . Positive differences between both conditions were also present beyond after fixation onset at some left frontal recording sites (e.g., F7 and F3), though less consistent. We statistically analyzed these effects using a cluster-based paired permutation t-test, which identifies significant clusters of adjacent data points (defined temporally and spatially).
@@
The results from the permutation test with the alternative hypothesis reveal one significant cluster () in the time interval from to (cf. Figure 3c). Although almost all electrodes are considered to be part of the significant cluster (at least, at one time point), the largest temporal extension of significant amplitude differences can be observed at midline frontal and fronto-central recording sites (i.e., Fz, FCz). Figure 3c displays the significant cluster across time and electrodes (surrounding black lines) over the amplitude difference in the grand average.
@@
These results show that the late frontal positivity is significantly greater for confusing code than for clean code, indicating that both conditions are processed differently starting at around after onset of the first fixation.